OpenStudy (anonymous):

Amy has total of 100 coins that are all dimes and quarters. The amount of money that equals is $21.40 Set up a system of equations and write the matrix equation that describes this situation. Let x = dimes and y = quarters.

OpenStudy (anonymous):

x + y = 100 .10x + .25y = 21.4

OpenStudy (anonymous):

Do you know how to use substitution?

OpenStudy (anonymous):

Or did you just need the equations?

OpenStudy (anonymous):

i wouldnt know how to set it up

OpenStudy (anonymous):

x + y = 100 is the same as y = -x + 100 So put the value of y into the other equation and solve for x: 0.1x + 0.25(-x + 100) = 21.4 0.1x + -0.25x + 25 = 21.4 combine like terms: -0.15x + 25 = 21.4 -25 -0.15x = -3.6 x = 24 Now plug the value of x into one of the original equations and solve for y: x + y = 100 24 + y = 100 -24 y = 76 So this means you have 24 dimes and 76 quarters. Hope I helped!